Find the value of the following expressions, when x = –1: 
 2x 2 – x – 2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that x = -1
Then, substitute the value of x in the question
= (2 × (-1)^2)–(-1)–2
= (2 × 1) + 1–2
=2 + 1–2
=3–2
=1

Thanks